# GUing: A Mobile GUI Search Engine using a Vision-Language Model

### v2

@@ 4. Evaluation Design
For the empirical evaluation of our search engine, we focus on answering the following research questions:
  • RQ1: How does GUing perform in text-to-GUI retrieval tasks compared to state-of-the-art approaches? As the CLIP model (Radford et al., 2021) was trained on 400 millions of image-caption pairs from the internet, it can also be applied on text-to-GUI retrieval task. An alternative method involves leveraging only the textual information: i.e. the text displayed on the GUI images, computing its semantic similarity with the input textual query. Thus, the question is whether the performance our proposed GUIClip model surpasses the CLIP model and text-only approaches.
@@
To answer these research questions, we designed two experiments: One (Exp1) consists of benchmarking three screenshot-caption datasets, and the other (Exp2) is based on a manual evaluation of the search engines.
4.1. Exp1: Evaluation of Model Performance on Three Datasets
@@ 4.1.4. Evaluation Metric
4.2. Exp2: Manual Evaluation of Search Engines
@@ 5. Evaluation Results
Table 4. Text-to-GUI retrieval performance on test set (Experiment 1)
@@
5.1. RQ1 Performance of GUIClip
@@
5.2. RQ2 Relevance of Search Results
Table 5. Relevance of the retrieved screenshots by three GUI search engines (Experiment 2 / Manual assessments)
@@ 5.2. RQ2 Relevance of Search Results
Figure 7. Top-10 screenshots returned by GUing for the query ”sleep tracking”.
@@
Top-10 screenshots returned by GUing for the query ”sleep tracking”.
@@
Figure 8. Top-10 screenshots returned by RaWi with query ”sleep tracking”
To understand the difference between the screenshots obtained using GUing and Rawi, we carried out additional analyses on each search engine with a few queries. For example, the screenshots returned by GUing and RaWi for the query ”sleeping track” are shown on Figure 8 and 8. Upon examination, it became apparent that there are considerable stylistic differences between the screenshots returned by the two search engines, mainly due to three factors:
@@
6. Exploring GUIClip as a Foundation Model for GUI Tasks
In the previous section, we evaluated the performance of GUIClip (as a main component of GUing) for the text-to-GUI retrieval task. In this section, we address the following research question:
RQ3: Does GUIClip also show a high performance for other GUI-related tasks? The CLIP model, as a foundation model, has been extensively applied across various vision-language tasks, consistently delivering promising results (Zhang et al., 2024). As GUIClip is a fine-tuned version specifically with GUI images, it has the potential to perform well for other GUI-related tasks. Particularly, the question is whether GUIClip can also outperform CLIP in other GUI-related tasks.
To explore this, we conducted two experiments: one focusing on GUI classification and one on sketch-to-GUI retrieval.
6.1. GUI Classification
GUI classification aims to categorize a screenshot as a whole under a specific label. It can be a fundamental activity for other GUI-related tasks, like GUI captioning (Leiva et al., 2023), the tagging of user screenshots (e.g. submitted in support requests) with certain requirements, components, or devices (Maalej et al., 2009; Roehm et al., 2013; Martens and Maalej, 2019a), as well as an early prototyping by a larger group and multiple screens (Pham et al., 2018). In this section, we explore the performance of GUIClip on GUI classification under zero-shot and linear-probe setting using the Enrico dataset (Leiva et al., 2020).
Figure 9. Overview of zero-shot and linear-probe classification on Enrico dataset with GUIClip (left side is adapted from (Radford et al., 2021)).
[…]
Overview of zero-shot and linear-probe classification on Enrico dataset with GUIClip (left side is adapted from (Radford et al., 2021)).
6.1.1. Model Architecture
6.1.2. Baselines
6.1.3. Dataset
6.1.4. Training Details of Linear-probe Classification
6.1.5. Evaluation Metric
Table 6. Classification accuracy on Enrico dataset
@@ 6.1.6. Results and Discussion
6.2. Sketch-to-GUI Retrieval
[…]
6.2.1. Model Architecture
6.2.2. Dataset
6.2.3. Baselines
  • Swire. Swire establishes itself as the pioneer sketch-to-GUI retrieval model (Huang et al., 2019). This model comprises two distinct VGG-A networks (Simonyan and Zisserman, 2015) that independently compute the embeddings for screenshots and sketches. The retrieval process is subsequently conducted by measuring the similarity between these embeddings. Given that Swire’s implementation is not publicly accessible, and since we were unable to get it from the authors, we took it upon ourselves to recreate it according to the details in the original paper.
6.2.4. Training Details of Fine-tuning
For the training of Swire, we tried to use the same configurations as specified in the original paper, including a learning rate of 0.01 and a batch size of 32. However, we noticed a significant shortfall in performance with these settings, wherein the top-10 recall stood at a mere 0.02. To address this, we made the decision to utilise a batch size of 64 and a learning rate of $1e^{-5}$. The original article does not specify the number of training epochs employed. Thus, we proceeded to train it for 100 epochs utilising the training set, until we observe convergence in the training loss.
6.2.5. Evaluation Metric
Table 7. Retrieval accuracy on the test set of Swire dataset
@@ 6.2.6. Results and Discussion
The performance of our Swire implementation, with a recall@1 of 0.027 and a recall@10 of 0.214, is not congruent with the metrics reported in the originating study by Huang et al. (Huang et al., 2019). The researchers reported a significantly higher recall values of 0.159 and 0.609, at ranks 1 and 10 respectively. This discrepancy could potentially be attributed to two factors: 1) Our evaluation used 486 pairings for testing, in contrast to the 276 pairings used in the original research. The size of the test set can dramatically influence retrieval performance outcomes; 2) The implementation details could also lead to differing results. It is important to note, however, that the original Swire model’s implementation is not available for replication. Nonetheless, even if we take into consideration the results from Huang et al. (Huang et al., 2019), our Dual GUIClip ViT model still demonstrates superior performance, with a score of 0.368 for recall@1 and 0.772 for recall@10.
@@ 7.2. Vision Language Models for Mobile GUI
Vision-language models combines both the vision and language modalities. They are trained on image-text pairs and can be applied to a range of tasks (Du et al., 2022; Zhang et al., 2024). Until this work, the main use case of VLMs in the mobile GUI domain is the GUI captioning. This refers to the task of generating a high-level summary of a screenshot to describe its contents and functionalities. The Screen2Words model (Wang et al., 2021), built on the foundations of the Transformer Encoder-Decoder architecture (Vaswani et al., 2017), uses multi-modal data, inclusive of the screenshot image, view hierarchy, and app description, to yield the screen caption. Contrarily, Spotlight, a vision-only approach employed for multiple tasks including widget captioning and screen captioning, possess architecture rooted in Vision Transformer (Dosovitskiy et al., 2021) and the T5 model (Raffel et al., 2020).
While some studies pertaining to GUI-related tasks may appear to employ VLMs, they do not in reality. For instance, XUI (Leiva et al., 2023) model does not fully address the language modality. It generates the screen caption with template-based natural language generation (NLG) engine. As discussed above, GUIGLE (Bernal-Cardenas et al., 2019) and RaWi (Kolthoff et al., 2023) introduced text-based GUI retrieval approaches by using the textual content exhibited in the screenshots or the metadata, instead of employing the vision information for the search. This may result in the omission of crucial vision information of the UI screen. To the best of our knowledge, no existing research employs VLMs for text-to-GUI retrieval.
Recently, two related VLMs, UIClip (Wu et al., 2024) and Ferret-UI (You et al., 2024), have been presented in parallel to our work. UIClip, like our model GUIClip, is a fine-tuned version the CLIP model. However, their purposes differ significantly. While GUIClip is designed for GUI retrieval (as app vendors describe what is implemented on the corresponding screen), UIClip is trained to assess the design quality and visual relevance of a UI. Ferret-UI, on the other hand, is a multimodal foundation model that goes beyond VLMs by supporting more input formats, such as points, boxes, and scribbles on screenshots. This model is capable of performing various tasks (e.g., widget classification, icon recognition, OCR) with flexible input formats, as well as grounding tasks (e.g., finding widgets, icons, or text, and listing widgets on a screen).
@@ 7.3. Mobile GUI Datasets
Large-scale mobile GUI datasets comprise a vast collection of screenshots and serve as a valuable resource for GUI retrieval. Rico (Deka et al., 2017; Liu et al., 2018) is one of the largest datasets in the literature. Deka et al. (Deka et al., 2017) combined crowdsourcing and automation to mine design and interaction data from Android apps at runtime. Their dataset exposes visual, textual, structural, and interactive design properties of more than 66k unique UI screens appx. 9.3k Android apps. Several other datasets (Leiva et al., 2020; Sunkara et al., 2022; Zaki and Abdallah, 2023) are built on Rico, including the Screen2Words dataset (Wang et al., 2021) used in this paper. As Rico was created seven years ago, some of its UIs might be outdated for contemporary apps as our evaluation suggests.
@@
Generally, more and more crowdsourced or automated GUI exploration methods are used to (semi)automatically explore UIs of iOS apps (Wu et al., 2023) or Android apps (Chen et al., 2019a, 2023; Zhang et al., 2023). While certainly efficient and scalable, automated UI exploration rather targets UI testing than GUI design inspiration. GUI datasets gathered through crowdsourced or automated exploration may omit important app features sine the access to certain UIs may necessitate app authorisation or initial configurations (Chen et al., 2023). In addition, these datasets are not curated by actual app designers and vendors. Recent studies showed that app vendors and app users (or the crowd) often use different vocabularies to describe app features (Haering et al., 2021a, b). The dataset introduced in our work ScreenRepo and SCapRepo aims at closing this gap. While we also used automated techniques for the data collection and cleaning, the screen-caption pairs in SCapRep are curated by various app vendors, describing what a screen is supposed to do and what problem for the user does it solve.
@@ 8.1. Implications for Software Practitioners
Our work has several potential implications for software practitioners, particularly for app designers and requirements engineers. Our search engine primarily enables app designers and developers to get inspirations on how their screens can look like, for instance in an early project phase, when only a list of app features is available. Particularly in small development teams, which often lack resources and expertise for GUI design and usability engineering (Bornoe and Stage, 2013), this can provide a substantial support. Despite the general trend with Generative AI for more code generation and automation, we think that it remains crucial in professional app development to have the designer (or developer) in the loop (Wei et al., 2024). GUing is such a tool that uses AI to speed-up design tasks while ensuring the crucial role of human creativity.
@@
Given the significant impact of GUI on user experience and retention (Chen et al., 2021), and as GUI designs and user preferences evolve rapidly, it is essential for app designers to continuously observe and learn from trends and best practices. GUing leverages app introduction images curated by actual domain and app vendors. Our pipeline for image processing and model training allows for a seamless addition of new images to the GUing repository, ensuring a high quality and up-to-dateness. This helps designers learn the latest UI design trends. When augmented with download, rating, and feedback trends, retrieved screens can help not only get inspiration but also estimate user trends (Martens and Maalej, 2019b).
@@ 8.2. Implication for Software Researchers
There has been extensive research on app store mining over the last decade, with the majority of studies focusing on app reviews and descriptions (Al-Subaihin et al., 2021; Martin et al., 2017; Maalej et al., 2024). Less work have delved into the analysis of other developer information shared in the stores, such as app introduction images (Chen et al., 2019b; Feng et al., 2022). Our work highlights the potential of this information, particularly as it is curated, usually exhibits a high quality (as created by professionals), and captures various modalities (image, text, videos, etc.). We think that researchers should focus more on investigating this information and combining it with crowd-data for building multi-modal foundation models for software engineering.
From the app introduction images, we developed and shared the SCapRepo dataset and subsequently the vision-language model GUIClip. The evaluation results confirms the power for these models to learn and combine modalities, not only for screenshot retrieval but also at least for GUI classification and sketch-to-GUI retrieval. We think that GUIClip and similar foundation models tuned on software design data have the potential to be adapted for other GUI-related tasks, including GUI captioning and GUI generation. As a foundational model, CLIP model itself is the basis to numerous models (Zhang et al., 2024), such as the image captioning model ClipCap (Mokady et al., 2021) and the image generation model Stable Diffusion (Rombach et al., 2022). By replacing the CLIP model with tuned, domain specific models such as GUIClip, it is possible to enhance these derivative models’ capabilities in GUI captioning and GUI generation. Researchers should investigate this in details, not only with regard to model performance (or accuracy) but also with regard to quality of output, risk of prediction faults, as well as design and engineering workflows to be adapted (Wei et al., 2024).

### v3

@@ 4. Evaluation Design
For the empirical evaluation of our search engine as well as our vision-language model, we focus on answering the following research questions:
  • RQ1: How does GUIClip perform in text-to-GUI retrieval tasks compared to state-of-the-art approaches? As the CLIP model (Radford et al., 2021) was trained on 400 millions of image-caption pairs from the internet, it can also be applied on text-to-GUI retrieval task. An alternative method involves leveraging only the textual information: i.e. the text displayed on the GUI images, computing its semantic similarity with the input textual query. Thus, the question is whether the performance our proposed GUIClip model surpasses the CLIP model and text-only approaches.
@@
  • RQ3: Does GUIClip also show a high performance in other GUI-related tasks? The CLIP model, as a foundation model, has been extensively applied across various vision-language tasks, consistently delivering promising results (Zhang et al., 2024). As GUIClip is a fine-tuned version specifically with GUI images, it has the potential to perform well for other GUI-related tasks. Particularly, the question is whether GUIClip can also outperform CLIP in other GUI-related tasks.
To answer these research questions, we designed four experiments. The first two experiments evaluate our approach on text-to-GUI retrieval: the first experiment (Exp1) addresses RQ1 by benchmarking three screenshot-caption datasets, while the second (Exp2) addresses RQ2 through a manual evaluation of search engines. The final two experiments focus on RQ3, assessing the performance of GUIClip on more GUI-related tasks: the third experiment (Exp3) focusing on GUI classification and the fourth (Exp4) on sketch-to-GUI retrieval.
4.1. Exp1: Evaluation of GUIClip for Text-to-GUI Retrieval
@@ 4.1.4. Evaluation Metric
4.2. Exp2: Manual Evaluation of GUing
@@ 4.2.3. Evaluation Metric
4.3. Exp3: Evaluation of GUIClip for GUI Classification
GUI classification aims to categorise a screenshot as a whole under a specific label. It can be a fundamental activity for other GUI-related tasks, like GUI captioning (Leiva et al., 2023), the tagging of user screenshots (e.g. submitted in support requests) with certain requirements, components, or devices (Maalej et al., 2009; Roehm et al., 2013; Martens and Maalej, 2019a), as well as an early prototyping by a larger group and multiple screens (Pham et al., 2018). In this section, we explore the performance of GUIClip on GUI classification under zero-shot and linear-probe setting using the Enrico dataset (Leiva et al., 2020).
Figure 7. Overview of zero-shot and linear-probe classification on Enrico dataset with GUIClip (left side adapted from Radfort et al. (Radford et al., 2021)).
[…]
Overview of zero-shot and linear-probe classification on Enrico dataset with GUIClip (left side adapted from Radford et al. (Radford et al., 2021)).
4.3.1. Model Architecture
4.3.2. Baselines
4.3.3. Dataset
4.3.4. Training Details of Linear-probe Classification
4.3.5. Evaluation Metric
4.4. Exp4: Evaluation of GUIClip for Sketch-to-GUI Retrieval
[…]
4.4.1. Model Architecture
4.4.2. Dataset
4.4.3. Baselines
  • Swire. Swire establishes itself as the pioneer sketch-to-GUI retrieval model (Huang et al., 2019). This model comprises two distinct VGG-A networks (Simonyan and Zisserman, 2015) that independently compute the embeddings for screenshots and sketches. The retrieval process is subsequently conducted by measuring the similarity between these embeddings. As the implementation of Swire is not publicly accessible and as we were unable to get it from the authors, we decided to recreate it according to the details in the original paper.
4.4.4. Training Details of Fine-tuning
For the training of Swire, we tried to use the same configurations as specified in the original paper, including a learning rate of 0.01 and a batch size of 32. However, we noticed a significant decline in performance with these settings, wherein the top-10 recall stood at a mere 0.02. To address this, we decided to use a batch size of 64 and a learning rate of $1e^{-5}$. As the original article does not specify the number of training epochs employed, we explored different values, ranging from 10 to 100 epochs in increments of 10. Our experiments indicates that 100 epochs produce the best results.
4.4.5. Evaluation Metric
@@ 5. Evaluation Results
Table 4. Text-to-GUI retrieval performance on test set (Experiment 1).
@@
5.1. RQ1: Performance of GUIClip in Text-to-GUI Retrieval
@@
5.2. RQ2: Relevance of Search Results
Table 5. Relevance of the retrieved screenshots by three GUI search engines (Experiment 2 / Manual assessments).
@@ 5.2. RQ2: Relevance of Search Results
Figure 9. Top-10 screenshots returned by GUing for the query “sleep tracking”.
@@
Top-10 screenshots returned by GUing for the query “sleep tracking”.
@@
Figure 10. Top-10 screenshots returned by RaWi for the query “sleep tracking”.
To understand the difference between the screenshots obtained using GUing and Rawi, we carried out additional analyses on each search engine with a few queries. For example, the screenshots returned by GUing and RaWi for the query ”sleeping track” are shown on Figure 10 and 10. Upon examination, it became apparent that there are considerable stylistic differences between the screenshots returned by the two search engines, mainly due to three factors:
@@
5.3. RQ3: Performance of GUIClip in Other GUI-related Tasks
5.3.1. GUI Classification
Table 6. Classification accuracy on Enrico dataset.
@@ 5.3.1. GUI Classification
5.3.2. Sketch-to-GUI Retrieval
Table 7. Retrieval accuracy on the test set of Swire dataset.
@@ 5.3.2. Sketch-to-GUI Retrieval
The performance of our Swire implementation, with a recall@1 of 0.027 and a recall@10 of 0.214, is not congruent with the metrics reported in the originating study by Huang et al. (Huang et al., 2019). The researchers reported a significantly higher recall values of 0.159 and 0.609, at ranks 1 and 10 respectively. This discrepancy could potentially be attributed to two factors: 1) Our evaluation used 486 pairings for testing, in contrast to the 276 pairings used in the original research. The size of the test set can dramatically influence retrieval performance outcomes; 2) The implementation details could also lead to differing results. It is important to note, however, that the original Swire model’s implementation is not available for replication. As reported in Section 4.4, we tried our best to not disadvantage Swire in our evaluation as the original implementation was not accessible. Nonetheless, even if we use the results from Huang et al. (Huang et al., 2019), our Dual GUIClip ViT model still demonstrates superior performance, with a score of 0.368 for recall@1 and 0.772 for recall@10.
@@ 6.2. Vision Language Models for Mobile GUI
Vision-language models combines both the vision and language modalities. They are trained on image-text pairs and can be applied to a range of tasks (Du et al., 2022; Zhang et al., 2024). Until this work, the main use case of vision-language models in the mobile GUI domain is the GUI captioning. This refers to the task of generating a high-level summary of a screenshot to describe its contents and functionalities. The Screen2Words model (Wang et al., 2021), built on the foundations of the Transformer Encoder-Decoder architecture (Vaswani et al., 2017), uses multimodal data, inclusive of the screenshot image, view hierarchy, and app description, to yield the screen caption. Contrarily, Spotlight, a vision-only approach employed for multiple tasks including widget captioning and screen captioning, possess architecture rooted in Vision Transformer (Dosovitskiy et al., 2021) and the T5 model (Raffel et al., 2020).
While some studies pertaining to GUI-related tasks may appear to employ vision-language models, they do not in reality. For instance, XUI (Leiva et al., 2023) model does not fully address the language modality. It generates the screen caption with template-based natural language generation (NLG) engine. As discussed above, GUIGLE (Bernal-Cardenas et al., 2019) and RaWi (Kolthoff et al., 2023) introduced text-based GUI retrieval approaches by using the textual content exhibited in the screenshots or the metadata, instead of employing the vision information for the search. This may result in the omission of crucial vision information of the UI screen. To the best of our knowledge, no existing research employs vision-language models for text-to-GUI retrieval.
Recently, two related vision-language models, UIClip (Wu et al., 2024) and Ferret-UI (You et al., 2024), have been presented in parallel to our work. UIClip, like our model GUIClip, is a fine-tuned version the CLIP model. However, their purposes differ significantly. While GUIClip is designed for GUI retrieval (as app vendors describe what is implemented on the corresponding screen), UIClip is trained to assess the design quality and visual relevance of a UI. Ferret-UI, on the other hand, is a multimodal foundation model that goes beyond vision-language models by supporting more input formats, such as points, boxes, and scribbles on screenshots. This model is capable of performing various tasks (e.g., widget classification, icon recognition, OCR) with flexible input formats, as well as grounding tasks (e.g., finding widgets, icons, or text, and listing widgets on a screen).
@@ 6.3. Mobile GUI Datasets
Large-scale mobile GUI datasets comprise a vast collection of screenshots and serve as a valuable resource for GUI retrieval. Rico (Deka et al., 2017; Liu et al., 2018) is one of the largest datasets in the literature. Deka et al. (Deka et al., 2017) combined crowdsourcing and automation to mine design and interaction data from Android apps at runtime. Their dataset exposes visual, textual, structural, and interactive design properties of more than 66k unique UI screens appx. 9.3k Android apps. Several other datasets (Leiva et al., 2020; Sunkara et al., 2022; Zaki and Abdallah, 2023) are built on Rico including the Screen2Words dataset (Wang et al., 2021) used in this paper. As Rico was created seven years ago, some of its UIs might be outdated for contemporary apps as our evaluation suggests.
@@
Generally, more and more crowdsourced or automated GUI exploration methods are used to (semi)automatically explore UIs of iOS apps (Wu et al., 2023) or Android apps (Chen et al., 2019a, 2023; Zhang et al., 2023). While certainly efficient and scalable, automated UI exploration rather targets UI testing than GUI design inspiration. GUI datasets gathered through crowdsourced or automated exploration may omit important app features since the access to certain UIs may necessitate app authorisation or initial configurations (Chen et al., 2023). In addition, these datasets are not curated by actual app designers and vendors. Recent studies showed that app vendors and app users (or the crowd) often use different vocabularies to describe app features (Haering et al., 2021a, b). The dataset introduced in our work ScreenRepo and SCapRepo aims at closing this gap. While we also used automated techniques for the data collection and cleaning, the screen-caption pairs in SCapRep are curated by various app vendors, describing what a screen is supposed to do and what problem for the user does it solve.
@@ 7.1. Implications for Software Practitioners
Our work has several potential implications for software practitioners, particularly for app designers and requirements engineers. Our search engine primarily enables app designers and developers to get inspirations on how their screens can look like, for instance in an early project phase, when only a list of app features is available (Wei et al., 2024b). Particularly in small development teams, which often lack resources and expertise for GUI design and usability engineering (Bornoe and Stage, 2013), this can provide a substantial support. Despite the general trend with Generative AI for more code generation and automation, we think that it remains crucial in professional app development to have the designer (or developer) in the loop (Wei et al., 2024a). GUing is such a tool that uses AI to speed-up design while ensuring the crucial role of human creativity and oversight.
@@
Given the significant impact of GUI on user experience and retention (Chen et al., 2021), and as GUI designs and user preferences evolve rapidly, it is essential for app designers to continuously observe and learn from trends and best practices. GUing leverages app introduction images curated by actual app vendors and experts. Our pipeline for image processing and model training allows for a seamless addition of new images to the GUing repository, ensuring a high quality and up-to-dateness. This helps designers learn the latest UI design trends. When augmented with download, rating, and feedback trends, retrieved screens can help not only get inspiration but also estimate user trends (Martens and Maalej, 2019b).
@@ 7.2. Implication for Software Researchers
There has been extensive research on app store mining over the last decade, with the majority of studies focusing on app reviews and descriptions (Al-Subaihin et al., 2021; Martin et al., 2017; Maalej et al., 2024). Less work have delved into the analysis of other developer information shared in the stores, such as app introduction images (Chen et al., 2019b; Feng et al., 2022). Our work highlights the potential of this information, particularly as it is curated, usually exhibits a high quality (as created by professionals), and captures various modalities (image, text, videos, binaries, etc.). We think that researchers should focus more on investigating this information and combining it with crowd-data for building multimodal foundation models for software engineering.
From the app introduction images, we developed and shared the SCapRepo dataset and subsequently the vision-language model GUIClip. The evaluation results confirms the power for these models to learn and combine modalities, not only for screenshot retrieval but also at least for GUI classification and sketch-to-GUI retrieval. We think that GUIClip and similar foundation models tuned on software design data have the potential to be adapted for other GUI-related tasks, including GUI captioning, GUI generation, or GUI testing. As a foundational model, CLIP model itself is the basis to numerous models (Zhang et al., 2024), such as the image captioning model ClipCap (Mokady et al., 2021) and the image generation model Stable Diffusion (Rombach et al., 2022). By replacing the CLIP model with tuned, domain specific models such as GUIClip, it is possible to enhance these derivative models’ capabilities for specific tasks. Researchers should investigate this in details, not only with regard to model performance (or accuracy) but also with regard to quality of output, risk of prediction faults, as well as design and engineering workflows to be adapted and integrated (Wei et al., 2024a; Maalej, 2009).